### v1

# A translation of the paper “ Presentation of some observations that could be made to shed light on Meteorology

A translation of the paper “ Presentation of some observations that
could be made to shed light on Meteorology
by Johann Heinrich Lambert (1771),
from the (old-French language) paper: “ Exposé de quelques
obfervations qu’on pourroit faire pour répandre
du jour fur la Météorologie
 ” in the revue
Memoires de l’Academie Royale des Sciences
et Belles-Lettres de Berlin.
Translated by Dr. Hab. Pascal Marquet
Possible contact at: pascalmarquet@yahoo.com
Web Google-sites: https://sites.google.com/view/pascal-marquet
ArXiv: https://arxiv.org/find/all/1/all:+AND+pascal+marquet/0/1/0/all/0/1
Research-Gate: https://www.researchgate.net/profile/Pascal-Marquet/research
Version-1 / July 17, 2025
—————————————————————————————————————– [Uncaptioned image]
[Uncaptioned image] —————————————————————————————————————–
Lambert propounded the ideology of observing periodic phenomena first, try to derive their rules and then gradually expand the theory. He expressed his purpose in meteorology in the present translated 1771 paper.
Note that Johann Heinrich Lambert (26 or 28 August 1728 – 25 September 1777) is the same polymath who also defined the azimuthal, cylindrical and conical equal area, as well as conformal conic, map projections still used in Meteorology to plot the weather maps: the so-called Lambert’s projections.
Note also that Lambert (1779) will derive in a next paper an absolute lower limit for the temperatures (at about 270270-270- 270°C), long after the first attempt by Guillaume Amontons (1703) (at about 240240-240- 240°C), but far before the definition by William Thomson (1848, next Lord Kelvin) of the so-called absolute Kelvin’s scale of temperature still used nowadays (at about 273273-273- 273°C).
——– The English translated 1771 paper ———–
It seems to me that to make Meteorology more scientific than it is, we should imitate the Astronomers who, without considering all the details, begin by establishing general laws and average movements. In this way they are able to take anomalies into account, and subject them to laws. In this way they finally managed to predict phenomena with an accuracy that inspires, even to the most ignorant, respect for Astronomy.
What is the same about meteorology? It is very certain that it has general laws, and that a large number of periodic phenomena enter into it. But these latter can hardly yet be guessed. The observations made so far, among which there is no connection, are of little importance. Many changes, especially in the gravity of air, have their causes in distant countries. The important point, therefore, would be to understand their sequence. To achieve this, neither great expense nor much preparation will be required. And if considerable sums have been spent on some astronomical observations, meteorology, which is of such close interest to the whole human race, would also deserve to be devoted to it. This is what it all comes down to.
It is known that great barometric variations occur at the same time, or almost at the same time, over a large area of country, and I am not mistaken in saying that the course of the barometer is, with a few slight differences, the same in Russia as in Portugal. However, these analogous variations extend more in longitude than in latitude. For if, for example, the barometer under the Pole varies by 3 inches ($88$ cm), it only varies by 2 inches ($5.55.5$ cm) in Paris, and we know that between the Tropics its variations do not go beyond 3 or 4 lines ($77$ to $99$ mm). 111 Note that the old French inches (of $2.7072.707$ cm) was separated into $1212$ lines (of $2.2562.256$ mm), with one French foot equal to $1212$ French inches or $32.48432.484$ cm, which is about 1/3.07813.0781/3.0781 / 3.078 of a meter. This is what we observed, but we do not know how it happens. Perhaps the shrinking of Zones and Climates is contributing to this. But observations made according to a certain method will establish them for us infinitely better and in more detail. If it were absolutely feasible, I would divide the surface of the Earth into 20 equal triangles, therefore in the shape of an isocahedron (see the Figs. 1), and at the center of each of these triangles meteorological observations would be made. In addition to these 20 Observers, 12 could be placed at the points of intersection of these triangles; and thus with 32 Observers, one would be able to keep a record of all the revolutions of the Atmosphere which would be subject to some general law. The whole weight of the atmosphere, its equilibrium, and the way in which it changes and re-establishes itself could be taken into account. These observers would be separated from each other by 37 to 41 degrees; and this distance would not be too great, seeing that the variations of the barometer, which depend on some general law, extend, if not further, at least just as far.
—————————————————————————————————————– Refer to caption                   Refer to caption —————————————————————————————————————–
Figure 1: Two modern 3D visions of isocahedron: On the left from https://www.aquaportail.com/dictionnaire/definition/11550/icosaedre; and on the right from https://fr-academic.com/dic.nsf/frwiki/833227
This project could be carried out, if one were willing to bear the expense. Nothing is easier than to determine these triangles on the globe. But one can also deviate more or less from this absolute regularity; and then the expense is reduced to an unimportant thing. Instead of placing observers at regular, but perhaps inaccessible, inconvenient, and dangerous points, it would be enough to choose places where trading nations have established colonies, where there are missionaries, in a word, where there are people who know how to count and write; for that is all that is needed for meteorological observations.
In passing over the globe I looked for the places, or at least the surroundings of the places which would form, or almost form, triangles such as I have just described (see the Figure 2). It is true that more than one Nation should be interested in arranging things so that in such and such a year the meteorological observations would begin at the same time, in accordance with the Instruction which would be drawn up in this regard.
—————————————————————————————————————– Refer to caption —————————————————————————————————————–
Figure 2: The figure inserted between the pages 64 and 65 of Lambert (1771) (from: https://pictures.abebooks.com/inventory/31862847446.jpg)
I believe, however, that the English nation could begin to set the tone, and there is no doubt that other trading nations would contribute to complete the whole system. For this purpose, it would first be sufficient to publish in the Gazettes the resolution that has been taken. And the Royal Society of Sciences of London, after having drawn up the Instruction and the plan for these observation forces, would publish this plan, and send it to the Academies established in other trading countries. And as in Europe there are many amateurs of meteorological observations, I believe that several of them would conform to it without difficulty, since at least for Europe the simultaneous and successive state of the Atmosphere would be known to the point of finding its more particular laws.
I do not know whether the Pilots, when recording the winds in their logs and even determining the strength of the wind by the speed of the ship and by the state of the sails and the drift, also record the weather; but nothing could be easier. And although a ship at sea is a mobile observatory, the large number of ships which are always on a voyage would not fail to make known the state of the Atmosphere in these intermediate regions where observers could not be found.
To avoid making these observations difficult, until the results are seen, a barometer and a thermometer may suffice at first. The scales will be either corresponding or at least intelligible. 222 For instance either Centigrade, Celsius, Réaumur or Fahrenheit well-known temperature scales, with known formulas to derive the ones from the others. The locations where observations will be made will be, as far as possible, at sea level, in order to have the weight of the entire atmosphere everywhere. The weather will be noted, as well as the winds and their different degrees of strength. As for the time, and to take into account the difference of the Meridians, one could take the noon of London. In this way one will observe the evening in the East Indies, and the morning in the West Indies, in order to have observations made at the same time. We could have copies of the registers sent to us every year.
Here is how I believe they could be arranged.
They can be divided into four parts.
The first part will look at barometric observations. Each month will be given a folio page. This page will have as many columns as there are places where observations were made. And next to each day, the height of the barometer will be marked, which in each column corresponds to that day. The columns will follow the order of the longitude of the places. As for the barometric heights, it will be sufficient to express them in lines (of $2.2562.256$ mm) and decimal parts of lines. One could even limit oneself to marking by how much they exceed the $300300$ duodecimal lines of the foot; this will be to narrow the columns and to make comparison easier.
The second part will be arranged in the same way with respect to the thermometer, which can be that of Fahrenheit, because its zero is low enough that the difference between hot (i.e. positive Fahrenheit values >17.8absent17.8>-17.8> - 17.8°C) and cold (i.e. negative Fahrenheit values) need only be marked very rarely.
The third part will be arranged in the same way with respect to the winds. One, two, three points would be added to the letters S, N, W, O, to mark one, two, three degrees, etc. of force. But with regard to winds which rise or cease suddenly, it would be good to mark there also the time when they began or ended.
Finally, the fourth part, which will be arranged in the same way, will offer the visible and sensible state of the air. As in each column there is a square cell for each day, this cell can be left blank for serene days. As for the other days, one could mark the symbols shown in the left part of the Fig. 3, and in this way one could also, by doubling these signs, mark the degree of strength and duration, with for example those in the right part of the Fig. 3.
It is clear that in this way the simultaneous and successive state of the air will be obvious, and that it will be seen as if at a glance.
Figure 3: Symbols written at the bottom of the page 63 of Lambert (1771).
In the comparison that can then be made between these observations, small changes will hardly come into play, since they ordinarily depend on accidental causes, which do not extend their effect very far. But the great variations will be of greater consequence, in that they extend further and have effects of a longer duration. It is especially to these (great variations) that we must have regard, in order to discover the general laws of the simultaneous and successive sequence of time, and in what manner its changes influence from one country to another, from one climate to another, and finally from one hemisphere to another.
Thus, for example, in Europe there are sometimes entire months which are fine or rainy. How far does this extend and from which country does the change then come? This is where we may have occasion to find general laws. For example, can the causes of these changes be found on the ocean, or do they always come from some dry land?
It is also for this reason that in Europe the North-East wind brings good weather, while the opposite South-West wind brings rain. Where does this begin and end? And why is it that sometimes, and especially in certain seasons, this does not happen, and that when the North or East wind brings rain, it is usually a very widespread and very long-lasting rain?
Moreover, despite all the small irregularities, there is something periodic in the variations of the weather, even if it is only the daily and annual variations of cold and heat. This must lead to other more or less periodic phenomena. Now, by considering them in this way, on a large scale and relative to the whole Atmosphere, there is no doubt that we will find general laws and different causes of the anomalies which are considerable. There are still a thousand questions which can be clarified by means of these kinds of registers. We will see, for example, which districts are covered by clouds, what difference there is according to the seasons, what path the clouds take, how far the strong and variable winds extend, where they originate, how they counterbalance each other, where this happens, what relation they have to the variation of the weight of the air, why is it that, while during the equinox a part of the atmosphere passes the Equator, this hardly causes the barometer to vary at all, whereas under the Pole it can vary by 3 inches? etc.
If, at least in the main locations, we could add to the observations indicated above those of the quantity of rain, the evaporation of the hygrometer, the declination and inclination of the compass, the system would become more complete. As for hygrometers, I have already shown in my Essay on Hygrometry (Mém. de l’Acad. 1769), and I will show it again on other occasions, that these instruments can be made to correspond. Very recent observations, made over a whole year in Silesia by the Baron de Felbiger, very worthy Prelate of the Abbey of Sagan, and in Berlin by myself, have shown me that the variations in humidity are very similar and very often equal in these two locations.